OpenStudy (surana):

Given b = -4 and h =-2, what is the equation of the graph if the parent function is y = /-x?

OpenStudy (surana):

\[y = \sqrt{x}\] is what it should look like.

OpenStudy (surana):

And possible answers are A \[y = \sqrt{4x+8}\] B \[y = \sqrt{4x-8}\] C \[y = \sqrt{-4x-8}\] D \[y = \sqrt{-4x+8}\]

OpenStudy (surana):

@misty1212 Could you help, please?

OpenStudy (anonymous):

i probably sound like a broken record, but there is no standard notation for "b" or "a" or "h" with a "parent function" this is something some math teacher (probably with a degree in math ed) made up that being said, if you can look in your book or notes or something, can you find an example with \(f(x)\) and \(b\) and \(h\)? then maybe we can figure it out, it can't be that hard

OpenStudy (surana):

Let me see about finding one.

OpenStudy (surana):

I think it's supposed to look like \[f(x) \sqrt{2x-4}\] in terms of what it should be looking like with the b and h presumaably in the area of the 2 and the 4.

OpenStudy (anonymous):

you need an example where they say \(f(x)=\) something , then \(b=\) some number and \(h=\) some number and see what they do

OpenStudy (surana):

Okay...I'll try to find one.

OpenStudy (surana):

I couldn't find one, but the thing itself should look like: \[y=a \sqrt{b(x-h)} +k \]

OpenStudy (surana):

I'm not sure if that'll help, though.

OpenStudy (surana):

Thoughts?

OpenStudy (surana):

I think it's either C or D.

OpenStudy (anonymous):

\[\sqrt{-4(x-(-2))}=\sqrt{-4(x+2)}=\sqrt{-4x-8}\]

OpenStudy (surana):

So it's C, then.

OpenStudy (surana):

Thank you.

OpenStudy (anonymous):

yw